BexleyBrooks:

Rewrite in simplest terms: minus, 9, left parenthesis, 8, n, plus, 4, p, right parenthesis, minus, 7, p, minus, 2, left parenthesis, minus, 10, p, minus, 9, n, right parenthesis−9(8n+4p)−7p−2(−10p−9n)

MakaylaChuck23:



BexleyBrooks:

@makaylachuck23 wrote:
Created with RaphaëlReply Using Drawing
Not really helpful but thats cute :)

MakaylaChuck23:

To simplify the expression − 9 ( 8 𝑛 + 4 𝑝 ) − 7 𝑝 − 2 ( − 10 𝑝 − 9 𝑛 ) −9(8n+4p)−7p−2(−10p−9n), follow these steps: Distribute − 9 −9 and − 2 −2: − 9 ( 8 𝑛 ) − 9 ( 4 𝑝 ) − 7 𝑝 + 2 ( 10 𝑝 ) + 2 ( 9 𝑛 ) −9(8n)−9(4p)−7p+2(10p)+2(9n) This gives: − 72 𝑛 − 36 𝑝 − 7 𝑝 + 20 𝑝 + 18 𝑛 −72n−36p−7p+20p+18n Combine like terms: For 𝑛 n: − 72 𝑛 + 18 𝑛 = − 54 𝑛 −72n+18n=−54n For 𝑝 p: − 36 𝑝 − 7 𝑝 + 20 𝑝 = − 23 𝑝 −36p−7p+20p=−23p So, the simplified expression is: − 54 𝑛 − 23 𝑝 −54n−23p